# Lennart Berg and Torben Wiedemann

### v1

@@ 4 Foldings
for all $i<j\in\{1,\ldots,6\}$, where $\lambda\in\mathcal{R}$. For each $\beta\in D_{6}$, we define a root group $\tilde{U}_{\beta}^{D}\coloneqq\tilde{\theta}_{\beta}^{D}(\mathcal{R})$ and we denote by $\operatorname{Ch}(D_{6},\mathcal{R})$ the group generated by all root groups. This is a Chevalley group of type $D_{6}$. Since $GH_{3}=\pi(D_{6})$ up to scaling and since the map $\pi\colon D_{6}\rightarrow\pi(D_{6})$ is a bijection (see 4.5), we may also define $\tilde{U}_{\beta}^{GH}\coloneqq\tilde{U}_{\beta^{\prime}}^{D}$ for all $\beta\in GH_{3}$ where $\beta^{\prime}$ is the unique root in $D_{6}$ with $\pi(\beta^{\prime})=\beta$. In other words, $\tilde{U}_{\alpha}^{D}=\tilde{U}_{\pi(\alpha)}^{GH}$ for all $\alpha\in D_{6}$.
Now let $\alpha\in H_{3}$ be arbitrary. We define a corresponding root group by
$\tilde{U}_{\alpha}^{H}\coloneqq\langle\tilde{U}_{\beta}^{D}\mid\beta\in D_{6}\cap(\phi\circ\pi)^{-1}(\alpha)\rangle=\langle\tilde{U}_{\beta}^{GH}\mid\beta\in GH_{3}\cap\phi^{-1}(\alpha)\rangle.$
Using the notation of 4.6, we have $D_{6}\cap(\phi\circ\pi)^{-1}(\alpha)=\{\beta_{1},\beta_{2}\}$ where $\beta_{1}\coloneqq\kappa^{\text{sh}}(\alpha)$ and $\beta_{2}\coloneqq\kappa^{\text{lo}}(\alpha)$. By 4.7, the groups $\tilde{U}_{\beta_{1}}^{D}$ and $\tilde{U}_{\beta_{2}}^{D}$ commute, so $\tilde{U}_{\alpha}^{H}=\tilde{U}_{\beta_{1}}^{D}\tilde{U}_{\beta_{2}}^{D}=\tilde{U}_{\beta_{2}}^{D}\tilde{U}_{\beta_{1}}^{D}$. Hence the root homomorphism
@@
$\tilde{\theta}_{\alpha}^{H}\colon\mathcal{R}\times\mathcal{R}\rightarrow\tilde{U}_{\alpha}^{H},(x,y)\mapsto\begin{cases}\tilde{\theta}_{\beta_{1}}^{D}(x)\tilde{\theta}_{\beta_{2}}^{D}(y)&\text{if }\alpha\notin\{\rho_{1},-\rho_{1}\},\\ \tilde{\theta}_{\beta_{1}}^{D}(x)\tilde{\theta}_{\beta_{2}}^{D}(-y)&\text{if }\alpha\in\{\rho_{1},-\rho_{1}\}\end{cases}$
is bijective. The reason to define $\tilde{\theta}_{\alpha_{13}}^{H}$ in a “twisted” way is of aesthetic nature: It ensures that the $A_{2}$-pair $(\rho_{1},\rho_{2})=(\alpha_{13},\alpha_{7})$ of simple roots satisfies the commutator relation
$\left[\tilde{\theta}_{\rho_{1}}^{H}(a,b),\tilde{\theta}_{\rho_{2}}^{H}(c,d)\right]=\tilde{\theta}_{\rho_{1}+\rho_{2}}^{H}(ac,bd)$
for $(a,b),(c,d)\in\mathcal{R}\times\mathcal{R}$ (without an additional sign in this relation).
@@
The construction of the root groups $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ in 4.10 applies to any $D_{6}$-graded group $(G,(\tilde{U}_{\alpha}^{D})_{\alpha\in D_{6}})$ in place of $\operatorname{Ch}(D_{6},\mathcal{R})$. By [Wie23, 5.7.14], any such group $G$ has a parametrisation $(\tilde{\theta}_{\alpha}^{D}\colon(\mathcal{R}^{\prime},+)\rightarrow\tilde{U}_{\alpha}^{D})_{\alpha\in D_{6}}$ by a commutative ring $\mathcal{R}^{\prime}$, a result which was first proven in [Shi93]. This allows us to define root isomorphisms $(\tilde{\theta}_{\alpha}^{H}\colon(\mathcal{R}^{\prime}\times\mathcal{R}^{\prime},+)\rightarrow\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ as in 4.10. The main result 4.18 of this section and its proof remain true in this generality except that, in order for $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ to satisfy Axiom 3.4 (iv), we have to require that $\tilde{U}_{-\alpha}^{D}\tilde{U}_{-\alpha^{\prime}}^{D}\cap\tilde{U}_{\Pi_{D}}^{D}=\{1_{G}\}$ for all $\alpha\in\Pi_{D}$ where $\alpha^{\prime}$ is the unique root in $\Pi_{D}\setminus\{\alpha\}$ with $\phi\pi(\alpha^{\prime})=\phi\pi(\alpha)$. This requirement is only slightly stronger than Axiom 3.4 (iv) for $(\tilde{U}_{\alpha}^{D})_{\alpha\in D_{6}}$, and always satisfied if $\tilde{U}_{\Pi_{D}}^{D}\cap\tilde{U}_{-\Pi_{D}}^{D}=\{1\}$.
Notation 4.12.
For the rest of this section, we fix the notation from 4.10.
Before we can prove that $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ is an $H_{3}$-grading of $\operatorname{Ch}(D_{6},R)$, we have to verify some properties of the chain of maps $D_{6}\rightarrow GH_{3}\rightarrow H_{3}$ and state (the mild generalisation of) a general result from [Wie23].
@@
Let $(i,j)\in\{(1,6),(3,5),(2,4)\}$, so that $r\coloneqq\sigma_{\delta_{i}}\sigma_{\delta_{j}}$ is an element of $R$ in 4.5. Then $\pi(\alpha^{r})=\pi(\alpha)^{r}=\pi(\alpha)^{\sigma(\pi(\delta_{i}))}=\pi(\alpha)^{\sigma(\pi(\delta_{j}))}$ for all $\alpha\in D_{6}$. Further, $\kappa^{\text{sh}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}=\kappa^{\text{sh}}(\gamma^{\sigma(\pi(\delta_{i}))})$ and $\kappa^{\text{lo}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}=\kappa^{\text{lo}}(\gamma^{\sigma(\pi(\delta_{i}))})$ for all $\gamma\in H_{3}$.
@@
Let $\alpha\in D_{6}$. Since $\pi$ is an orthogonal projection, we have $\pi(\alpha^{r})=\pi(\alpha)^{r}$. The assertion that $\pi(\alpha)^{r}=\pi(\alpha)^{\sigma(\pi(\delta_{i}))}=\pi(\alpha)^{\sigma(\pi(\delta_{j}))}$ is a straightforward computation, see 2.11 for details. Now let $\gamma\in H_{3}$. Then for $\beta\in\{\kappa^{\text{lo}},\kappa^{\text{sh}}\}$,
$\phi\circ\pi\big{(}\beta(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}\big{)}$ $=\phi\big{(}\pi(\beta(\gamma))^{\sigma(\pi(\delta_{i}))}\big{)}=\big{(}\phi\circ\pi(\beta(\gamma)))^{\sigma(\pi(\delta_{i}))}=\gamma^{\sigma(\pi(\delta_{i}))}.$
Further, for all $\gamma\in H_{3}$,
$\pi\big{(}\kappa^{\text{lo}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}\big{)}$ $=\pi\big{(}\kappa^{\text{lo}}(\gamma)\big{)}^{\sigma(\pi(\delta_{i}))}=\tau\cdot\pi\big{(}\kappa^{\text{sh}}(\gamma)\big{)}^{\sigma(\pi(\delta_{i}))}$
$=\tau\cdot\pi\big{(}\kappa^{\text{sh}}(\gamma)^{\sigma(\delta_{i})\sigma(\delta_{j})}\big{)}.$
The remaining assertion follows. ∎
@@
For all invertible $r\in\mathcal{R}$ and all $\beta\in D_{6}$, it is well-known that $w^{D}_{\beta}(r)\coloneqq\tilde{\theta}_{-\beta}^{D}(-r^{-1})\tilde{\theta}_{\alpha}^{D}(r)\tilde{\theta}_{-\beta}^{D}(-r^{-1})$ is a $\beta$-Weyl element in the $D_{6}$-grading $(\tilde{U}_{\gamma}^{D})_{\gamma\in D_{6}}$ (see [Ste16, Lemma 20 (b)] and [Wie23, 3.3.6]). Now let $\alpha\in H_{3}$ and let $\beta_{1}\coloneqq\kappa^{\text{sh}}(\alpha),\beta_{2}\coloneqq\kappa^{\text{lo}}(\alpha)\in D_{6}$ be as in 4.6. For invertible $r,s\in\mathcal{R}$, we put
$w^{H}_{\alpha}(r,s)\coloneqq w^{D}_{\beta_{1}}(r)w^{D}_{\beta_{2}}(s)=\tilde{\theta}_{-\alpha}^{H}(-r^{-1},-s^{-1})\tilde{\theta}_{\alpha}^{H}(r,s)\tilde{\theta}_{-\alpha}^{H}(-r^{-1},-s^{-1}).$
The proof of [Wie23, 2.6.6] applies verbatim in this situation, except that references to [Wie23, 1.4.10] must be replaced by 4.13. We conclude that $w^{H}_{\alpha}(r,s)$ is an $\alpha$-Weyl element with respect to $(\tilde{U}_{\gamma}^{H})_{\gamma\in H_{3}}$. Alternatively, it is a straightforward computation to verify this fact.
Lemma 4.15.
Let $\alpha,\beta\in GH_{3}$ (or $\alpha,\beta\in H_{3}$). Then there exists a positive system $\Pi$ in $D_{6}$ which contains $D_{6}\cap\pi^{-1}(\{\alpha,\beta\})$ (or $D_{6}\cap(\phi\pi)^{-1}(\{\alpha,\beta\})$).
@@
Let $\alpha,\beta\in GH_{3}$. Then $\alpha,\beta$ lie in a common positive system in $GH_{3}$, which must be conjugate to the system $\Pi_{GH}$ from 4.8. Hence there exist $\zeta_{1},\ldots,\zeta_{k}\in\{\rho_{1},\rho_{2},\rho_{3}\}=\Delta_{H}$ such that
$\alpha,\beta\in(\Pi_{GH})^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}=(\pi(\Pi_{D}))^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}.$
By 4.13, there exists $w\in\operatorname{Weyl}(D_{6})$ such that $(\pi(\Pi_{D}))^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}=\pi((\Pi_{D})^{w})$. Then $\pi(-(\Pi_{D})^{w})=-(\pi(\Pi_{D}))^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})}$ as well, so $\pi^{-1}((\Pi_{GH})^{\sigma(\zeta_{1})\cdots\sigma(\zeta_{k})})=(\Pi_{D})^{w}$. Thus $\Pi\coloneqq(\Pi_{D})^{w}$ is a positive system in $D_{6}$ with the desired property. The same choice of $\Pi$ works for $\alpha^{\prime}\coloneqq\phi(\alpha),\beta^{\prime}\coloneqq\phi(\beta)\in H_{3}$. ∎
Lemma 4.16.
Let $\bar{\pi}\colon\Phi\rightarrow\Phi^{\prime}$ be a surjective map between root systems $\Phi,\Phi^{\prime}$ such that $\Phi$ is reduced. Assume that for all $\alpha^{\prime},\beta^{\prime}\in\Phi^{\prime}$, there exists a positive system $\Pi$ in $\Phi$ which contains $\bar{\pi}^{-1}(\{\alpha^{\prime},\beta^{\prime}\})$. Let $G$ be a group which has (crystallographic) $\Phi$-commutator relations with root groups $(\bar{V}_{\alpha})_{\alpha\in\Phi}$ and put $\bar{U}_{\alpha^{\prime}}\coloneqq\bar{V}_{\bar{\pi}^{-1}(\alpha^{\prime})}$ for all $\alpha^{\prime}\in\Phi^{\prime}$. If, in the notation of 2.6,
$\bar{\pi}(\mathord{]A,B[})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}\qquad\text{for all }A,B\subseteq\Phi,$
then $(\bar{U}_{\alpha^{\prime}})_{\alpha^{\prime}\in\Phi^{\prime}}$ satisfies Axiom 3.3 (i). If the commutator relations of $G$ are crystallographic and
$\bar{\pi}(\mathord{]A,B[}_{\text{cry}})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}_{\text{cry}}\qquad\text{for all }A,B\subseteq\Phi,$
then $(\bar{U}_{\alpha^{\prime}})_{\alpha^{\prime}\in\Phi^{\prime}}$ satisfies the crystallographic version of Axiom 3.3 (i).
Proof.
This follows from the same arguments as in [Wie23, 2.1.29]. ∎

### v2

@@ 4 Foldings
for all $i<j\in\{1,\ldots,6\}$, where $\lambda\in\mathcal{R}$. For each $\beta\in D_{6}$, we define a root group $\tilde{U}_{\beta}^{D}\coloneqq\tilde{\theta}_{\beta}^{D,\text{n}}(\mathcal{R})$ and we denote by $\operatorname{Ch}(D_{6},\mathcal{R})$ the group generated by all root groups. This is a (simply connected) Chevalley group of type $D_{6}$, and hence a $D_{6}$-graded group. It is equipped with a family $(\tilde{\theta}_{\alpha}^{D,\text{n}})_{\alpha\in D_{6}}$ of root isomorphisms.
Example 4.23.
Let $\mathcal{R}$ be a commutative ring, let $L$ be a complex Lie algebra of type $E_{8}$ and let $(x_{\alpha})_{\alpha\in E_{8}}\cup(h_{i})_{1\leq i\leq 8}$ be a Chevalley basis of $L$ in the sense of [Hum78, 25.2]. Then by [Hum78, 25.5], $L(\mathbb{Z})\coloneqq\langle x_{\alpha},h_{i}\mid\alpha\in E_{8},1\leq i\leq 8\rangle_{\mathbb{Z}}$ is invariant under $\operatorname{ad}_{x_{\alpha}}^{n}/n!$ for all $\alpha\in E_{8}$ and $n\in\mathbb{N}_{0}$. Further, $\operatorname{ad}_{x_{\alpha}}$ is nilpotent for all $\alpha\in E_{8}$. Hence for all $\alpha\in E_{8}$, we may define a homomorphism
@@
$\tilde{\theta}_{\alpha}^{E,\text{n}}\colon(\mathcal{R},+)\rightarrow\operatorname{GL}_{\mathcal{R}}\big{(}L(\mathbb{Z})\otimes\mathcal{R}\big{)},\lambda\mapsto\sum_{n=0}^{\infty}\frac{\operatorname{ad}_{x_{\alpha}}^{n}}{n!}\otimes\lambda^{n}\operatorname{id}_{\mathcal{R}}$
and a corresponding root group $\tilde{U}_{\alpha}^{E}\coloneqq\tilde{\theta}_{\alpha}^{E,\text{n}}(\mathcal{R})$, which acts on $L(\mathbb{Z})\otimes\mathcal{R}$ on the left. (For details, see [VP96, Section 3].) The group generated by $(\tilde{U}_{\alpha}^{E})_{\alpha\in E_{8}}$ is an (adjoint) Chevalley group of type $E_{8}$, which we denote by $\operatorname{Ch}(E_{8},\mathcal{R})$. It is equipped with an $E_{8}$-grading $(\tilde{U}_{\alpha}^{E})_{\alpha\in E_{8}}$ and a family $(\tilde{\theta}_{\alpha}^{E,\text{n}})_{\alpha\in E_{8}}$ of root isomorphisms. Note that it is a matrix group of degree $\dim_{\mathbb{C}}L=\lvert E_{8}\rvert+\operatorname{rank}(E_{8})=248$.
We now show that $(U_{\beta}^{H})_{\beta\in H_{\ell}}$ is an $H_{\ell}$-grading of $G$. At first, we prove that there exist Weyl elements with respect to $(U_{\beta}^{H})_{\beta\in H_{\ell}}$.
Lemma 4.24.
Let $\beta\in H_{\ell}$ and write $\pi^{-1}(\beta)=\{\alpha_{1},\alpha_{2}\}$ where $\alpha_{1},\alpha_{2}\in X$. For each $i\in\{1,2\}$, let $w_{\alpha_{i}}$ be an $\alpha_{i}$-Weyl element in $(G,(U_{\gamma}^{X})_{\gamma\in X})$. Then $w_{\beta}\coloneqq w_{\alpha_{1}}w_{\alpha_{2}}=w_{\alpha_{2}}w_{\alpha_{1}}$ is a $\beta$-Weyl element in $(G,(U_{\delta}^{H})_{\delta\in H_{\ell}})$.
Proof.
By (the arguments in) 4.20, $\langle U_{\alpha_{1}}^{X},U_{-\alpha_{1}}^{X}\rangle$ commutes with $\langle U_{\alpha_{2}}^{X},U_{-\alpha_{2}}^{X}\rangle$. Hence
$w_{\beta}\in U_{-\alpha_{1}}^{X}U_{-\alpha_{2}}^{X}U_{\alpha_{1}}^{X}U_{\alpha_{2}}^{X}U_{-\alpha_{1}}^{X}U_{-\alpha_{2}}^{X}=U_{-\beta}^{H}U_{\beta}^{H}U_{-\beta}^{H}.$
Now let $\delta\in H_{\ell}$ be arbitrary and write $\pi^{-1}(\delta)=\{\gamma_{1},\gamma_{2}\}$ for $\gamma_{1},\gamma_{2}\in X$. Then
$(U_{\delta}^{H})^{w_{\beta}}$ $=(U_{\gamma_{1}}^{X}U_{\gamma_{2}}^{X})^{w_{\alpha_{1}}w_{\alpha_{2}}}=U_{\gamma_{1}^{\sigma(\alpha_{1})\sigma(\alpha_{2})}}^{X}U_{\gamma_{2}^{\sigma(\alpha_{1})\sigma(\alpha_{2})}}^{X}=U_{\gamma_{1}^{u(\sigma_{\beta})}}^{X}U_{\gamma_{2}^{u(\sigma_{\beta})}}^{X}$
by 4.17. Since $\{\gamma_{1}^{u(\sigma_{\beta})},\gamma_{2}^{u(\sigma_{\beta})}\}=\pi^{-1}(\delta)^{u(\sigma_{\beta})}=\pi^{-1}(\delta^{\sigma_{\beta}})$ by 2.14, we infer that $(U_{\delta}^{H})^{w_{\beta}}=U_{\delta^{\sigma(\beta)}}^{H}$. Hence $w_{\beta}$ is indeed a $\beta$-Weyl element. ∎
While 4.24 assures the existence of Weyl elements, we want to go further and compute explicit formulas for the actions of certain Weyl elements in the Chevalley groups considered previously. This boils down to computing certain structure constants, or rather, signs. It is part of our main result 9.33 that every $H_{\ell}$-graded group (for $\ell\geq 3$) can be coordinatised in such a way that the same structural signs as for foldings of Chevalley groups are satisfied. However, the structural signs that we would obtain from the families $(\tilde{\theta}_{\alpha}^{D,\text{n}})_{\alpha\in D_{6}}$ and $(\tilde{\theta}_{\alpha}^{E,\text{n}})_{\alpha\in E_{8}}$ in 4.22 and 4.23 do not behave as we would like (see 4.26 for details). Hence we will twist these families in the way described as follows, and then compute the desired signs in 4.27.
Definition 4.25.
Let $T$ be a subset of a positive system of $X$, let $\mathcal{R}$ be an abelian group and let $(\theta_{\alpha}^{X,\text{n}}\colon\mathcal{R}\rightarrow U_{\alpha}^{X})_{\alpha\in X}$ be a family of root isomorphisms. The $T$-twist of $(\theta_{\alpha}^{X,\text{n}})_{\alpha\in X}$ is the family $(\theta_{\alpha}^{X})_{\alpha\in X}$ defined by
$\theta_{\alpha}^{X}\colon\mathcal{R}\rightarrow U_{\alpha}^{X},\lambda\mapsto\begin{cases}\theta_{\alpha}^{X,\text{n}}(-\lambda)&\text{if }\alpha\in T\cup(-T),\\ \theta_{\alpha}^{X,\text{n}}(\lambda)&\text{if }\alpha\notin T\cup(-T).\end{cases}$
@@
For the group $\operatorname{Ch}(D_{6},\mathcal{R})$ in 4.22, we will choose
$T(D_{6})\coloneqq\{\delta_{6}=e_{5}+e_{6}\}.$
This choice will ensure that the formula for the commutator relation of the pair $(\tilde{U}_{\rho_{1}}^{H},\tilde{U}_{\rho_{2}}^{H})$ has a certain form. See 4.30 for details.
For the group $\operatorname{Ch}(E_{8},\mathcal{R})$ in 4.23, the situation is more complicated because, even without twisting, different choices of the Chevalley basis result in different structure constants of the Chevalley group. For our computation of explicit signs, we will use the Chevalley basis that is provided by the GAP [GAP24] command SimpleLieAlgebra("E", 8, Rationals) and the $T(E_{8})$-twist defined by
$T(E_{8})\coloneqq\left\{\begin{gathered}[0,0,0,1,1,1,0,0],[0,1,1,1,1,1,0,0],[0,0,1,2,1,1,0,0],\\ [0,0,0,0,0,1,0,0],[1,1,1,1,1,1,0,0],[0,0,1,1,1,1,0,0],\\ [1,1,1,2,1,1,0,0],[0,1,1,2,1,1,0,0],[0,1,2,2,1,1,0,0],\\ [1,1,2,2,1,1,0,0],[1,2,2,2,1,1,0,0]\end{gathered}\right\}.$
Here we use Notation 2.10. This choice ensures that the canonical $D_{6}$-graded subgroup of $\operatorname{Ch}(E_{8},\mathcal{R})$ has the same “structure constants” as the $D_{6}$-graded group in 4.22. For details, see 4.27 and 4.30.
In the following, we will write $(\tilde{\theta}_{\alpha}^{D})_{\alpha\in D_{6}}$ and $(\tilde{\theta}_{\alpha}^{E})_{\alpha\in E_{8}}$ for the twists of the families $(\tilde{\theta}_{\alpha}^{D,\text{n}})_{\alpha\in D_{6}}$ and $(\tilde{\theta}_{\alpha}^{E,\text{n}})_{\alpha\in E_{8}}$ (from 4.22 and 4.23) by the sets specified above. Further, we will write $(\tilde{\theta}_{\beta}^{H})_{\beta\in H_{\ell}}$ for the corresponding folding in the sense of 4.21.
Remark 4.27.
Let $\mathcal{R}$ be a commutative ring and let $X=E_{8}$. We consider the group $\operatorname{Ch}(E_{8},\mathcal{R})$ with the $E_{8}$-grading $(\tilde{U}_{\alpha}^{E})_{\alpha\in E_{8}}$ from 4.23 and the families $(\tilde{\theta}_{\alpha}^{E})_{\alpha\in E_{8}}$ and $(\tilde{\theta}_{\beta}^{H})_{\beta\in H_{4}}$ of root isomorphisms from 4.26. It is a well-known fact about Chevalley groups (see [Ste16, Lemma 20 (b)] and [Wie24, 3.3.6] for details) that for all invertible $r\in\mathcal{R}$ and all $\alpha\in E_{8}$, the element
$w_{\alpha}(r)\coloneqq\tilde{\theta}_{-\alpha}^{E}(-r^{-1})\tilde{\theta}_{\alpha}^{E}(r)\tilde{\theta}_{-\alpha}^{E}(-r^{-1})$
is an $\alpha$-Weyl element in $(G,(\tilde{U}_{\alpha}^{E})_{\alpha\in E_{8}})$. Hence it follows from 4.24 that for all invertible $r,s\in\mathcal{R}$ and all $\alpha_{1},\alpha_{2}\in E_{8}$ with $\tilde{\pi}(\alpha_{2})=\tau\tilde{\pi}(\alpha_{1})$ and $\beta\coloneqq\pi(\alpha_{1})=\pi(\alpha_{2})$, the element
$w_{\beta}(r,s)\coloneqq$ $\mathrel{}w_{\alpha_{1}}(r)w_{\alpha_{2}}(s)=w_{\alpha_{2}}(s)w_{\alpha_{1}}(r)$
$=$ $\mathrel{}\tilde{\theta}_{-\beta}^{H}(-r^{-1},-s^{-1})\tilde{\theta}_{\beta}^{H}(r,s)\tilde{\theta}_{-\beta}^{H}(-r^{-1},-s^{-1})$
is a $\beta$-Weyl element in $(G,(\tilde{U}_{\gamma}^{H})_{\gamma\in H_{\ell}})$. Put $w_{\rho}\coloneqq w_{\rho}(1_{\mathcal{R}},1_{\mathcal{R}})$ for all $\rho\in\{\rho_{0},\rho_{1},\rho_{2},\rho_{3}\}$. A straightforward but long matrix computation (see 2.18) shows that for all $\beta\in H_{4}$ and all $\rho\in\{\rho_{0},\rho_{1},\rho_{2},\rho_{3}\}$, there exist signs $\epsilon_{\beta,\delta},\bar{\epsilon}_{\beta,\delta}\in\{\pm 1\}$ such that
$\tilde{\theta}_{\beta}^{H}(r,s)^{w_{\rho}}=\tilde{\theta}_{\beta^{\sigma(\rho)}}^{H}(\epsilon_{\beta,\delta}r,\bar{\epsilon}_{\beta,\delta}s)$ (4.28)
for all $r,s\in\mathcal{R}$. If $2_{\mathcal{R}}\neq 0_{\mathcal{R}}$, then the signs $\epsilon_{\beta,\delta},\bar{\epsilon}_{\beta,\delta}$ are uniquely determined by (4.28) and they do not depend on $\mathcal{R}$. If, on the other hand, $2_{\mathcal{R}}=0_{\mathcal{R}}$, then $\epsilon_{\beta,\delta},\bar{\epsilon}_{\beta,\delta}$ may be chosen arbitrarily. We define
$\eta\colon H_{4}\times\{\rho_{0},\rho_{1},\rho_{2},\rho_{3}\}\rightarrow\{\pm 1\}\times\{\pm 1\},(\beta,\delta)\mapsto\eta_{\beta,\delta}=\eta(\beta,\delta)\coloneqq(\epsilon_{\beta,\delta},\bar{\epsilon}_{\beta,\delta})$
where $\epsilon_{\beta,\delta},\bar{\epsilon}_{\beta,\delta}$ denote the unique choice of signs that satisfies (4.28) for all commutative rings $\mathcal{R}$. The map $\eta$ is called the standard parity map for $H_{4}$. It will be studied in more detail in Section 6. For all $\beta\in H_{4}$ and $\delta\in\{\rho_{0},\rho_{1},\rho_{2},\rho_{3}\}$, the value of $\eta(\beta,\delta)$ can be computed in a straightforward way, and this yields that $\eta(-\beta,\delta)=\eta(\beta,\delta)$ (see 2.18). All values of $\eta(\beta,\delta)$ for positive $\beta$ are given in Figures 12 and 13 in the appendix. Note that, while $\eta$ describes important structure constants of $G$, it can be defined without reference to $G$ by simply listing all values of $\eta$, as in the aforementioned tables.
We can apply similar computations to the Chevalley group of type $D_{6}$ from 4.22 to obtain another parity map $\eta^{\prime}\colon H_{3}\times\{\rho_{1},\rho_{2},\rho_{3}\}\rightarrow\{\pm 1\}\times\{\pm 1\}$. The set $T(E_{8})$ in 4.26 has been chosen precisely in a way to ensure that $\eta(\beta,\delta)=\eta^{\prime}(\beta,\delta)$ for all $\delta\in\{\rho_{1},\rho_{2},\rho_{3}\}$ and all $\beta\in H_{4}\cap\langle\rho_{1},\rho_{2},\rho_{3}\rangle_{\mathbb{R}}$. In other words, $\eta^{\prime}$ is the restriction of $\eta$ to $H_{3}$. See 2.18 for a verification of this fact. For the reader’s convenience, the values of $\eta^{\prime}$ are given in Figure 4. In the following, we will simply write $\eta$ for $\eta^{\prime}$.
$[0,0,1]$ $(1,1)$ $(-1,-1)$ $(-1,-1)$
$[0,\tau,1]$ $(1,1)$ $(1,-1)$ $(1,1)$
$[0,\tau,\tau]$ $(-1,-1)$ $(1,1)$ $(1,-1)$
$[\tau,\tau,1]$ $(-1,1)$ $(1,1)$ $(-1,-1)$
$[0,1,\tau]$ $(-1,1)$ $(1,-1)$ $(-1,-1)$
$[\tau,\tau,\tau]$ $(1,-1)$ $(-1,-1)$ $(-1,1)$
$[0,1,0]$ $(-1,-1)$ $(-1,-1)$ $(1,-1)$
$[1,1,\tau]$ $(1,-1)$ $(1,1)$ $(-1,1)$
$[\tau,\tau^{2},\tau]$ $(-1,1)$ $(1,1)$ $(-1,-1)$
$[1,1,0]$ $(1,1)$ $(-1,-1)$ $(1,1)$
$[1,\tau^{2},\tau]$ $(-1,-1)$ $(-1,1)$ $(1,-1)$
$[\tau,\tau^{2},\tau^{2}]$ $(1,1)$ $(-1,-1)$ $(1,1)$
$[1,0,0]$ $(-1,-1)$ $(1,1)$ $(1,1)$
$[1,\tau^{2},\tau^{2}]$ $(1,-1)$ $(1,1)$ $(-1,1)$
$[\tau,2\tau,\tau^{2}]$ $(1,1)$ $(-1,1)$ $(1,1)$
Figure 4: Values of the parity map $\eta^{\prime}\colon H_{3}\times\{\rho_{1},\rho_{2},\rho_{3}\}\rightarrow\{\pm 1\}\times\{\pm 1\}$ in 4.27. The values $\eta_{\alpha,\delta}$ for negative roots $\alpha$ are given by the formula $\eta_{-\alpha,\delta}=\eta_{\alpha,\delta}$.
Recall the definition of commutation maps in 3.17. In the situation of the following 4.29, we have $M=\mathcal{R}\times\mathcal{R}$.
@@
Let $\mathcal{R}$ be a commutative ring, let $X\in\{D_{6},E_{8}\}$ and consider the group $\operatorname{Ch}(X,\mathcal{R})$ from 4.22 or 4.23. Further, let $(\tilde{\theta}_{\beta}^{H})_{\beta\in H_{\ell}}$ be the family of root isomorphisms from 4.26. Put $\alpha\coloneqq\rho_{2}$, $\varepsilon\coloneqq\rho_{3}$ and denote the corresponding $H_{2}$-quintuple by $(\alpha,\beta,\gamma,\delta,\varepsilon)$. Then for all non-proportional non-adjacent roots $\zeta,\xi\in E_{2}^{+}(\Delta_{H})$ (as defined in 2.8) and all $\rho\in\mathord{]\zeta,\xi[}$, the commutation map $\psi_{\zeta,\xi}^{\rho}$ is given by Figure 5.
@@
This is a straightforward (matrix) computation. See 2.18 for details. ∎
[…]
Figure 5: The commutator relations in the folding of a Chevalley group. See 4.29 for details.
@@
Similar to the behaviour of the parity map in 4.27, we observe that the commutator relations in Figure 5 on the canonical $H_{3}$-subsystem of $H_{4}$ hold for the group from 4.22 as well as for the group from 4.23. Again, this fact is due to our choice of $T(E_{8})$ in 4.26. Further, the inclusion of $\delta_{1}$ in $T(D_{6})$ and of $\delta_{6}$ in $T(E_{8})$ ensures that the formula for $\psi_{\rho_{1},\rho_{2}}$ does not involve signs.
Remark 4.31.
The commutator relations of the $X$-grading of $\operatorname{Ch}(X,\mathcal{R})$, which we will need in 10.8, are easier to state: For all roots $\xi,\zeta\in X$ for which $\xi+\zeta\in X$, there exists a sign $c_{\xi,\zeta}\in\{\pm 1\}$ such that
$[\tilde{\theta}_{\xi}^{X}(r),\tilde{\theta}_{\zeta}^{X}(s)]=\tilde{\theta}_{\xi+\zeta}^{X}(c_{\xi,\zeta}rs)$
for all $r,s\in\mathcal{R}$. See [VP96, (9.4)] for details, but note that [VP96] uses the non-twisted family $(\theta_{\xi}^{X,\text{n}})_{\xi\in X}$ of root isomorphisms.
Proposition 4.32.
Let $(G,(U_{\alpha}^{X})_{\alpha\in X})$ be any $X$-graded group. Assume that for all positive systems $\Pi$ in $X$ and all $\alpha_{1},\alpha_{2}\in\Pi$ with $\pi(-\alpha_{1})=\pi(-\alpha_{2})$, we have $U_{\Pi}^{X}\cap U_{\alpha_{1}}^{X}U_{\alpha_{2}}^{X}=\{1_{G}\}$. Then the folding $(U_{\beta}^{H})_{\beta\in H_{\ell}}$ defined in 4.19 is an $H_{\ell}$-grading of $G$.
@@
Since every root group in $(U_{\alpha}^{X})_{\alpha\in X}$ is contained in a root group in $(U_{\beta}^{H})_{\beta\in H_{\ell}}$, the groups $(U_{\beta}^{H})_{\beta\in H_{\ell}}$ generate $G$. Further, $(G,(U_{\beta}^{H})_{\beta\in H_{\ell}})$ has Weyl elements by 4.24.
Denote by $\Pi_{X}$ and $\Pi_{H}$ the positive systems in $X$ and $H_{\ell}$ corresponding to the root bases $\Delta_{X}$ and $\Delta_{H}$, respectively. Since $\pi(\Delta_{X})=\Delta_{H}$ by 4.8 a, we have $\pi(\Pi_{X})=\Pi_{H}$ and thus $U_{\Pi_{X}}^{X}=U_{\Pi_{H}}^{H}$. Hence our additional assumption on $U_{\Pi}^{X}\cap U_{-\alpha_{1}}^{X}U_{-\alpha_{2}}^{X}$ says precisely that $(U_{\beta}^{H})_{\beta\in H_{\ell}}$ satisfies Axiom 3.4 (iv).
It remains to show that $G$ has $H_{\ell}$-commutator relations with root groups $(U_{\beta}^{H})_{\beta\in H_{\ell}}$. For the special case of the Chevalley groups in 4.22 and 4.23, this is known by the matrix computations in 4.29. For arbitrary $X$-graded groups, this is more technical and will be proven in A.9 a in the appendix. ∎